What percent did Sarah score incorrectly if her score was 48 out of 64?
Answer by jim_thompson5910(35256)

She scored 48 correct answers.
She got 64 - 48 = 16 questions wrong.

The percent incorrect is
16/64 = 1/4 = 0.25 = 25%

Final Answer: 25%